# A localisation phase transition for the catalytic branching random walk

### v1

@@ Abstract
We show the existence of a phase transition between a localisation and a non-localisation regime for a branching random walk with a catalyst at the origin. More precisely, we consider a continuous-time branching random walk that jumps at rate one, with simple random walk jumps on $\mathbb{Z}^{d}$, and that branches (with binary branching) at rate $\lambda>0$ everywhere, except at the origin, where it branches at rate $\lambda_{0}>\lambda$. We show that, if $\lambda_{0}$ is large enough, then the occupation measure of the branching random walk localises (i.e. converges almost surely without spatial renormalisation), whereas, if $\lambda_{0}$ is close enough to $\lambda$, then localisation cannot occur, at least not in a strong sense. The case $\lambda=0$ (when branching only occurs at the origin) has been extensively studied in the literature and a transition between localisation and non-localisation was also exhibited in this case. Strikingly, the transition that we observe, conjecture, and partially prove in this paper occurs at the same threshold as in the case $\lambda=0$.
@@ 1 Introduction and main results
In this paper, we consider the following continuous-time branching random walk on $\mathbb{Z}^{d}$ ($d\geq 1$): at time zero, there is one particle alive in the system, and its position is the origin. Each particle carries two independent Poisson clocks: the “jump”-clock that rings at rate $1$, and the “branch”-clock that rings at rate $\lambda_{0}$ when the particle is at the origin and at rate $\lambda$ everywhere else. When the jump-clock of a particle rings, then this particle moves to a neighbouring site chosen uniformly at random among the $2d$ neighbouring sites. When the branch-clock of a particle rings, it gives birth to a new particle at the same location. This model is a variant of the so-called “catalytic branching random walk”, which corresponds to taking $\lambda=0$, and which has been extensively studied in the literature (we give a literature review on this model in Section 1.4).
@@
$\Pi_{t}=\sum_{i=1}^{N(t)}\delta_{X_{i}(t)},$
with $N(t)$ the number of particles alive at time $t$, and $(X_{i}(t))_{1\leq i\leq N(t)}$ their respective positions. We assume that $\lambda_{0}\geq\lambda>0$.
@@ 1.1 Results
If $\lambda_{0}>2d-1+2d\lambda$, then, there exists a deterministic measure $\nu$ on $\mathbb{Z}^{d}$ such that, almost surely as $t\uparrow\infty$,
$\widehat{\Pi}_{t}:=\frac{\Pi_{t}}{N(t)}\to\nu,$
@@
The strength of this result is that it gives almost sure convergence of the occupation measure of the branching random walk. In comparison, existing results in the case $\lambda=0$, such as in [DR13], only give convergence of all moments (see Section 1.4 for a more detailed discussion and more references on the case $\lambda=0$). Another notable difference is that we use a random normalisation, namely we divide the empirical distribution of particles by $N(t)$, and our limit is a deterministic measure. This contrasts for instance with limiting results in the case $\lambda=0$, such as in [DR13] or [Wat67] in a continuous setting, where the normalisation is deterministic and the limiting measure is multiplied by a random factor. We believe that our results (Theorem 1.1 and Lemma 2.1 below) may imply convergence of the occupation measure normalised by a deterministic factor. Proving this would require to estimate the speed of convergence in Theorem 1.1, which we have not tried to do so far.
The following results indicate that the condition $\lambda_{0}>2d-1+2d\lambda$ may not be optimal: i.e. we believe that localisation should occur for smaller values of $\lambda_{0}$. First we show that any candidate for the limiting measure $\nu$ satisfies some balance equations:
@@
In fact, (1.2) is redundant. Indeed, by summing over $x\neq 0$ the left and right hand sides of the equalities in (1.1), it is readily seen that any probability measure satisfying (1.1) also satisfies (1.2). Finally, our last result identifies a phase transition for the existence of solutions to the balance equations (1.1) and (1.2), which we conjecture to coincide with the transition between a localisation and a delocalisation regimes, see Conjecture 1.4 below.
@@
In other words, we prove that, $(i)$ if $\lambda_{0}>2d-1+2d\lambda$, then the occupation measure of the branching random walk “localises” (see Theorem 1.1), and $(ii)$ if $0\leq\lambda_{0}-\lambda\leq\gamma_{d}$, then localisation cannot occur in the strong sense of $(i)$ (see Theorem 1.3$(i)$). We stress that the value of the threshold for the transition observed in Theorem 1.3, namely the constant $\gamma_{d}$, is exactly the same as in the case $\lambda=0$, see below. However, while the phase transition is elementary to identify in the latter case, using a direct comparison between the total number of particles ever visiting the origin and a standard Galton-Watson process, it seems less immediate to use such a comparison when $\lambda$ is positive. Actually, it may even seem counter-intuitive that the threshold for the phase transition between localisation and delocalisation holds at the same critical value (for the parameter $\lambda_{0}-\lambda$), no matter the value of $\lambda$. Indeed, one could expect that increasing $\lambda$ would favour the natural diffusive behaviour of the process, while our results tend to show that, on the contrary, it does not play any role in this respect.
@@ 1.2 Discussion of the proofs (and plan of the paper)
The proof of localisation (see Section 4), i.e. Theorem 1.1, relies on proving that the occupation measure taken at the times when either a branching or a reproduction event occurs is a measure-valued Pólya process (mvpp). Mvpps were introduced in [MM17] and [BT16] (also see [BT17] were the model was first defined on a particular example) as a generalisation of Pólya urns to infinitely-many colours. Mailler and Marckert [MM17] proved convergence in probability of a large class of mvpps under an assumption of balance (in terms of urns, the balance assumption requires that the total number of balls added at any time-step in the urn is constant and deterministic). They also required that the replacement rule is deterministic, but this assumption was later removed by Janson [Jan19]. To prove Theorem 1.1, we use convergence results proved by Mailler and Villemonais [MV20], which apply without the balance assumption, for random replacements, and hold almost surely. Besides showing that our model is an mvpp, our main contribution is to check that our mvpp satisfies the assumptions of [MV20]; to do this, we need to show that a continuous-time sub-Markovian process admits a quasi-stationary distribution, and this is done using results of Champagnat and Villemonais [CV23]. Interestingly, this quasi-stationary distribution is the limit $\nu$ in Theorem 1.1, and our Proposition 1.2 provides a characterisation of this measure for $\lambda_{0}$ large enough.
@@
The proof of Proposition 1.2 (see Section 2) relies on (1) finding a good approximation for the total number of particles alive at large times $t$ (see Lemma 2.1 - this is done by finding a martingale and proving that it converges in $L^{2}$), which allows to replace the random normalisation of the occupation measure by a deterministic one (see Lemma 2.2), and then (2) writing some forward evolution equation for the occupation measure and studying their equilibrium.
@@
Finally, the proof of Theorem 1.3 (see Section 3) is done by studying the existence (or lack of thereof) solutions to the balance equations of Proposition 1.2. To do this, we use the optional stopping theorem applied to a martingale that naturally arises from the balance equations.
1.3 Conjectures
As briefly mentioned before, when $\lambda_{0}=\lambda$, it is well-known that, almost surely as $t\uparrow\infty$
$\widehat{\Pi}_{t}\big{(}\cdot\sqrt{t}\big{)}\to\mathcal{N}(0,1),$
for the topology of weak convergence on the space of probability measures on $\mathbb{R}$. This implies in particular that, for all $B\subset\mathbb{Z}$, almost surely as $t\uparrow\infty$,
$\widehat{\Pi}_{t}(B)\to 0.$
We conjecture that the phase transition is sharp, i.e. that this “diffusive” behaviour stays true for $\lambda_{0}$ larger than $\lambda$, up to a critical threshold after which the occupation measure localises. In view of Theorem 1.3, we conjecture that this threshold equals $\lambda+\gamma_{d}$:
Conjecture 1.4.
  1. (i)
    For all $\lambda\leq\lambda_{0}\leq\lambda+\gamma_{d}$, for all $B\subset\mathbb{Z}$, almost surely as $t\uparrow\infty$,
    $\widehat{\Pi}_{t}(B)\to 0.$
  2. (ii)
    For all $\lambda_{0}>\lambda+\gamma_{d}$, almost surely as $t\uparrow\infty$,
    $\widehat{\Pi}_{t}\to\nu,$
    for the topology of weak convergence on the space of probability measures on $\mathbb{Z}^{d}$, where $\nu$ is the unique probability measure satisfying (1.1).
In the case $\lambda\leq\lambda_{0}\leq\lambda+\gamma_{d}$ (assuming $\gamma_{d}>0$), one could even wonder if there exists a renormalisation function $(u_{t})_{t\geq 0}$ and a limiting measure $\mu$ such that $\widehat{\Pi}_{t}(\cdot\,u_{t})\to\mu$. As discussed above, we know from the literature that, if $\lambda=\lambda_{0}$, then such a result holds with $u_{t}=\sqrt{t}$ and $\mu$ the standard Gaussian. But the question of how do $u_{t}$ and $\mu$ change when $\lambda_{0}$ increases remains open.
@@ 1.4 Discussion of the literature
To compare the aforementioned results with ours, we now summarise the results of, e.g. [DR13]. For all $y\in\mathbb{Z}^{d}$, $t\geq 0$, and $k\geq 1$, we let $M^{k}(t,y)$ be the $k$-th moment of the number of particles at position $y$ at time $t$, and $M^{k}(t)$ be the $k$-th moment of the total number of particles in the system at time $t$. Note that Döring and Roberts assume that branching at zero happens at rate 1, but allow the offspring distribution to be more general than binary fission. However, as explained above, one can interpret $m$, the average number of offspring, as the rate of branching at 0, i.e. $\lambda_{0}=m-1$. Indeed, to create $m$ offspring per unit of time from binary fissions, we need to have $m-1$ binary fissions per unit of time. Roberts and Döring’s results can thus be interpreted as:
  • if $\lambda_{0}<\gamma_{d}$, then, for all $y\in\mathbb{Z}^{d}$,
    $\lim_{t\uparrow\infty}M^{1}(t,x)\in(0,\infty)\quad\text{ and }\quad\lim_{t\uparrow\infty}M^{1}(t,x,y)=0,$
@@
  • If $\lambda_{0}>\gamma_{d}$, then, there exists a positive constant $\rho>0$ such that, for all $y\in\mathbb{Z}^{d}$,
    $\lim_{t\uparrow\infty}\mathrm{e}^{-k\rho t}M^{1}(t,y)\in(0,\infty)\quad\text{ and }\quad\lim_{t\uparrow\infty}\mathrm{e}^{-k\rho t}M^{1}(t)\in(0,\infty),$
@@
Note that the critical case, when $\lambda_{0}=\gamma_{d}$, is also studied in [DR13], and the growth rate of the number of particles in this case is sub-exponential (see also [VT11]).
@@
Generalisations of the catalytic branching random walk: Since its introduction, the catalytic branching random walk (with branching only at the origin) has been generalised in several ways. One direction of generalisation is to add more catalysts (sites at which branching happens): instead of having one catalyst at the origin, one can have $N$ catalysts at positions $x_{1},\ldots,x_{N}$ (see, e.g. [AB00]). Another direction is to allow the catalysts to move following their own Markovian motion and branching only happens for particles at the same position as one of the catalysts, with a rate that may depend on the number of catalysts there (see, e.g. [GdH06, GH06, KS03, CGM11]). Finally let us mention the model of inhomogeneous branching random walk, where the branching rate depends on the position of the particle, as e.g. in [BH14b, GKW99].

### v2

# Catalytic branching random walk

@@ Abstract
We show the existence of a phase transition between a localisation and a non-localisation regime for a branching random walk with a catalyst at the origin. More precisely, we consider a continuous-time branching random walk that jumps at rate one, with simple random walk jumps on $\mathbb{Z}^{d}$, and that branches (with binary branching) at rate $\lambda>0$ everywhere, except at the origin, where it branches at rate $\lambda_{0}>\lambda$. We show that, if $\lambda_{0}$ is large enough, then the occupation measure of the branching random walk localises (i.e. when normalised by the total number of particles, it converges almost surely without spatial renormalisation), whereas, if $\lambda_{0}$ is close enough to $\lambda$, then the occupation measure delocalises, in the sense that the proportion of particles in any finite given set converges almost surely to zero. The case $\lambda=0$ (when branching only occurs at the origin) has been extensively studied in the literature and a transition between localisation and non-localisation was also exhibited in this case. Interestingly, the transition that we observe, conjecture, and partially prove in this paper occurs at the same threshold as in the case $\lambda=0$. One of the strengths of our result is that, in the localisation regime, we are able to prove convergence of the occupation measure, whilst existing results in the case $\lambda=0$ give convergence of moments instead.
@@ 1 Introduction and main results
Branching random walks are a simple probabilistic model for populations that move through space. In the classical continuous time model, particles all branch at the same rate and between branching events, move according to some Markov process (eg. a simple random walk), independently of each other. A natural generalisation of this model is to allow the branching rate to depend on the position of the particle: from an application’s point of view, one can think of some sites being more favourable to a growth in population than others (for example if there is more food available at those sites). In some models such as the parabolic Anderson model, in which the branching rates are sampled in an i.i.d. way, some sites are sometimes so attractive compared to others that the population localises at these sites or near these sites. In this paper, we show that localisation can occur in a very simple model where the branching rate is the same everywhere, except at the origin, where it is higher.
We consider the following continuous-time branching random walk on $\mathbb{Z}^{d}$ ($d\geq 1$): at time zero, there is one particle alive in the system, and its position is the origin. Each particle carries two independent Poisson clocks: the “jump”-clock that rings at rate $1$, and the “branch”-clock that rings at rate $\lambda_{0}$ when the particle is at the origin and at rate $\lambda$ everywhere else. When the jump-clock of a particle rings, then this particle moves to a neighbouring site chosen uniformly at random among the $2d$ neighbouring sites. When the branch-clock of a particle rings, it gives birth to a new particle at the same location. This model is a variant of the so-called “catalytic branching random walk”, which corresponds to taking $\lambda=0$, and which has been extensively studied in the literature (we give a literature review on this model in Section 1.3).
@@
$\Pi_{t}=\sum_{i=1}^{N_{t}}\delta_{X_{i}(t)},$
with $N_{t}$ the number of particles alive at time $t$, and $(X_{i}(t))_{1\leq i\leq N_{t}}$ their respective positions. We assume in the whole paper that $\lambda_{0}\geq\lambda>0$ (in fact most of our results assume $\lambda_{0}>\lambda>0$, but we also occasionally discuss the case $\lambda_{0}=\lambda$).
@@ 1.1 Results
If $\lambda_{0}>2d-1+2d\lambda$, then there exists a deterministic probability measure $\nu$ on $\mathbb{Z}^{d}$ such that, almost surely as $t\uparrow\infty$,
$\widehat{\Pi}_{t}:=\frac{\Pi_{t}}{N_{t}}\to\nu,$
@@
The strength of this result is that it gives almost sure convergence of the occupation measure of the branching random walk. In comparison, existing results in the case $\lambda=0$, such as in [ABY98, DR13], only give convergence of all moments (see Section 1.3 for a more detailed discussion and more references on the case $\lambda=0$). Another notable difference is that we use a random normalisation, namely we divide the empirical distribution of particles by $N_{t}$, and our limit is a deterministic measure. This contrasts for instance with limiting results in the case $\lambda=0$, such as in Döring and Roberts [DR13] or Watanabe [Wat67] in a continuous setting, where the normalisation is deterministic and the limiting measure is multiplied by a random factor. We believe that our results (Theorem 1.1 and Lemma 3.1 below) may imply convergence of $\mathrm{e}^{-(\lambda+(\lambda_{0}-\lambda)\nu_{0})t}\Pi_{t}$. Proving this would require to estimate the speed of convergence in Theorem 1.1, which we have not tried to do so far.
We believe that the condition $\lambda_{0}>2d-1+2d\lambda$ is not optimal. To support that claim we first state the following proposition:
@@
In fact, (1.2) is redundant. Indeed, by summing over $x\neq 0$ the left and right hand sides of the equalities in (1.1), it is readily seen that any probability measure satisfying (1.1) also satisfies (1.2). Our next result identifies a phase transition for the existence of solutions to the balance equations (1.1) and (1.2), which we conjecture to coincide with the transition between a localisation and a delocalisation regimes.
@@
In other words, we prove that, (a) if $\lambda_{0}>2d-1+2d\lambda$, then the occupation measure of the branching random walk “localises” (see Theorem 1.1), and (b) if $0\leq\lambda_{0}-\lambda\leq\gamma_{d}$, then localisation cannot occur in the strong sense of (a) (see Theorem 1.3(i)(i)). We stress that the value of the threshold for the transition observed in Theorem 1.3, namely the constant $\gamma_{d}$, is exactly the same as in the case $\lambda=0$, see below. However, while the phase transition is elementary to identify in the latter case, using a direct comparison between the total number of particles ever visiting the origin and a standard Galton-Watson process (see Remark Remark below), it seems less immediate to use such a comparison when $\lambda$ is positive. Actually, it may even seem counter-intuitive that the threshold for the phase transition between localisation and delocalisation holds at the same critical value (for the parameter $\lambda_{0}-\lambda$), no matter the value of $\lambda$. Indeed, one could expect that increasing $\lambda$ would favour the natural diffusive behaviour of the process, while our results tend to show that, on the contrary, it does not play any role in this respect.
Our last main result shows a weak form of localisation, as soon as $\lambda_{0}>\lambda+1$, in the sense that almost surely at all large times, a positive proportion of all particles stands at the origin. It also shows that when $\lambda_{0}<\lambda+\gamma_{d}$, almost surely the proportion of particles that stand at the origin (or in any given finite set) converges to zero as time goes to infinity.
From now on, in a slight abuse of notation, we write $\widehat{\Pi}_{t}(x)$ instead of $\widehat{\Pi}_{t}(\{x\})$, for all $x\in\mathbb{Z}^{d}$.
Theorem 1.4.
  1. (i)
    If $\lambda_{0}>1+\lambda$, then, almost surely,
    $\liminf_{t\to\infty}\ \widehat{\Pi}_{t}(0)\geq\frac{\lambda_{0}-\lambda-1}{\lambda_{0}-\lambda}.$
  2. (ii)
    If $\lambda<\lambda_{0}<\lambda+\gamma_{d}$, then for any finite $B\subset\mathbb{Z}^{d}$, almost surely,
    $\lim_{t\to\infty}\widehat{\Pi}_{t}(B)=0.$
In addition to being interesting in itself, the first part of this theorem is used as a preliminary step in the proof of Theorem 1.1, see Remark Remark. Concerning the second part, we note that when $\lambda_{0}=\lambda$, it is well-known111 We were unable to find a reference where this is done in our exact setting, but one can apply standard martingale techniques as in [Big92] to get this result. that, almost surely as $t\uparrow\infty$,
$\widehat{\Pi}_{t}\big{(}\cdot\sqrt{t}\big{)}\to\mathcal{N}(0,\text{Id}),$
for the topology of weak convergence on the space of probability measures on $\mathbb{R}^{d}$, where Id stands for the identity matrix. Based on first moment estimates, we conjecture that such a strong result should persist in the whole subcritical phase, i.e. for any $\lambda\leq\lambda_{0}<\lambda+\gamma_{d}$.
Remark.
Several of the assumptions we make on the catalytic branching random walk are, we believe, not necessary (and are not generally made in the literature on the case $\lambda=0$). For example, the assumption that, at time zero, there is a unique particle and this particle sits at the origin could be replaced by having a finite number of particles alive at time zero, each of them at any site in $\mathbb{Z}^{d}$. This is because of the irreducibility of the process. Also, we believe that the binary branching could be replaced by any supercritical branching mechanism (even with a random number of offspring). Finally, the particles could move according to a Markov jump process instead of a continuous time simple symmetric random walk, as assumed here. Our proofs should be robust under all these modifications, although the proof of Theorem 1.1 might become more intricate, in particular in the case of a more general underlying Markov jump process. This is the reason why we decided to keep our framework as simple as possible.
@@ 1.2 Discussion of the proofs (and plan of the paper)
The proofs of Theorem 1.4 (i) and (ii) (see Section 2) rely on two different ideas. For the proof of the weak localisation result, we first show that the number of particles at the origin grows exponentially fast almost surely. For this we couple our catalytic branching random walk with the classical branching random walk (corresponding to the case $\lambda=\lambda_{0}$) and use a second moment method. We then use the excess of the branching rate at the origin and Borel-Cantelli type arguments, to bootstrap this estimate, and show that the proportion of particles at the origin remains bounded from below over time by a positive constant. The proof of the weak delocalisation result relies on a first moment method and on a direct link between the first moment in the case $\lambda_{0}>\lambda$ and the first moment in the case when branching only occurs at the origin, at rate $\lambda_{0}-\lambda$, see Proposition 2.5.
@@
The proof of Proposition 1.2 (see Section 3) relies on (1) finding a good approximation for the total number of particles alive at large times $t$ (see Lemma 3.1 - this is done by finding a martingale and proving that it converges in $L^{2}$), which allows to replace the random normalisation of the occupation measure by a deterministic one (see Lemma 3.2), and then (2) writing a forward evolution equation for the occupation measure and studying their equilibrium.
@@
The proof of Theorem 1.3 (see Section 4) is done by studying the existence of solutions (or lack of thereof) to the balance equations of Proposition 1.2. To do this, we use the optional stopping theorem applied to a martingale that naturally arises from the balance equations.
The proof of localisation (see Section 5), i.e. Theorem 1.1, relies on proving that the occupation measure taken at the times when either a branching or a jumping event occurs is a measure-valued Pólya process (mvpp). Mvpps were introduced in Bandyopadhyay and Thacker [BT16] and Mailler and Marckert [MM17] (see also [BT17] where the model was first defined on a particular example) as a generalisation of Pólya urns to infinitely-many colours. Mailler and Marckert [MM17] proved convergence in probability of a large class of mvpps under an assumption of balance (in terms of urns, the balance assumption requires that the total number of balls added at any time-step in the urn is constant and deterministic). They also required that the replacement rule is deterministic, but this assumption was later removed by Janson [Jan19]. To prove Theorem 1.1, we use convergence results proved by Mailler and Villemonais [MV20], which apply without the balance assumption, for random replacements, and hold almost surely. Besides showing that our model is an mvpp, our main contribution is to check that our mvpp satisfies the assumptions of [MV20]; to do this, we need to show that a continuous-time sub-Markovian process admits a quasi-stationary distribution, and this is done using results of Champagnat and Villemonais [CV23]. Interestingly, this quasi-stationary distribution is the limit $\nu$ in Theorem 1.1, and our Proposition 1.2 provides a characterisation of this measure for $\lambda_{0}$ large enough.
As already discussed, we believe that the condition $\lambda_{0}>2d-1+2d\lambda$ is far from optimal (and conjecture that there is localisation as soon as $\lambda_{0}>\lambda+\gamma_{d}$). The reason for this is that the proof of Theorem 1.1 relies on [MV20, Theorem 1] and [CV23, Theorem 5.1], both of which are of the form “sufficient assumptions imply result” as opposed to “necessary and sufficient assumptions imply result”. To improve our condition on $\lambda_{0}$, one would have to come up with an ad-hoc proof of these two theorems for the catalytic branching random walk, or come up with a completely different proof that does not involve MVPPs.
@@ 1.3 Discussion of the literature
To compare the aforementioned results with ours, we now summarise the results of, e.g. [ABY98] (see also [DR13] for similar results in a slightly different model). For all $y\in\mathbb{Z}^{d}$, $t\geq 0$, and $k\geq 1$, we let $M^{k}(t,y)$ be the $k$-th moment of the number of particles at position $y$ at time $t$, and $M^{k}(t)$ be the $k$-th moment of the total number of particles in the system at time $t$. Albeverio, Bogachev and Yarovaya [ABY98] prove that:
  • If $\lambda_{0}<\gamma_{d}$, then, there exist positive constants $C,(C(y))_{y\in\mathbb{Z}^{d}}$ such that, for all $y\in\mathbb{Z}^{d}$, as $t\to+\infty$,
    $M^{1}(t,y)=(C(y)+o(1))t^{-\nicefrac{{d}}{{2}}}\quad\text{ and }\quad M^{1}(t)=C+o(1),$ (1.3)
@@
  • If $\lambda_{0}>\gamma_{d}$, then, there exist positive constants $\rho>0$, $C$, $(C(y))_{y\in\mathbb{Z}^{d}}$ such that, for all $y\in\mathbb{Z}^{d}$, for all $k\geq 1$, as $t\to+\infty$,
    $M^{k}(t,y)=(C(y)+o(1))\mathrm{e}^{k\rho t}\quad\text{ and }\quad M^{k}(t)=(C+o(1))\mathrm{e}^{k\rho t},$
@@
Note that the critical case, when $\lambda_{0}=\gamma_{d}$, is also studied in [ABY98], and the growth rate of the number of particles in this case is sub-exponential (see also Vatutin and Topchii [VT11]).
We believe that, for most of our results, the assumption that $\lambda$ is positive could be dropped. In particular, when $\lambda=0$, the convergence result of Theorem 1.1 should hold almost surely on the event that the number of particles grows to infinity. The main challenge would be to prove an analogue of Lemma 2.1 below, showing that on the event of non-local extinction, the number of particles at the origin grows almost surely exponentially fast. While it could very well be that our proof of Lemma 2.1 could be adapted to this case, using in particular the asymptotics of the first two moments of $\Pi_{t}(0)$ (proved in [ABY98]), we have not tried to pursue in this direction.
@@
Generalisations of the catalytic branching random walk: Since its introduction, the catalytic branching random walk (with branching only at the origin) has been generalised in several ways. One direction of generalisation is to add more catalysts (sites at which branching happens): instead of having one catalyst at the origin, one can have $N$ catalysts at positions $x_{1},\ldots,x_{N}$ (see, e.g. Albeverio and Bogachev [AB00]). Another direction is to allow the catalysts to move following their own Markovian motion and branching only happens for particles at the same position as one of the catalysts, with a rate that may depend on the number of catalysts there (see, e.g. [GH06, KS03, CGM11]). Finally let us mention the model of inhomogeneous branching random walk, where the branching rate depends on the position of the particle, as e.g. in [BH14b, GKW99].
Multi-type Galton-Watson processes: A natural discrete-time version of our process is the following. Start with one particle at the origin say, at time 0. Then given two probability measures $\mu_{0}$ and $\mu$, at each unit of time, particles alive at that time give birth to a number of offspring distributed respectively as $\mu_{0}$ if at the origin or as $\mu$ if not at the origin, and die immediately after that, whilst new born particles immediately jump at random to a nearest neighbor. Interpreting the spatial positions of particles as their types, we can also view this process as a multi-type Galton-Watson process. There are now a number of deep results regarding the asymptotic behavior of the empirical distribution of types for such processes, see in particular [AD25] and [BBC25], and references therein. It would be interesting to investigate if for such process, a sharp phase transition could be derived, depending on the means of $\mu_{0}$ and $\mu$.